40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 191 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  8377   Fri Mar 29 19:58:24 2013 Gabriele, JenneFrogsLSCAnalog whitening filter of REFL55 not switching

Quote:

We discovered that the analog whitening filter of the REFL55_I board is not switching when we operate the button on the user interface. We checked with the Stanford analyzer that the transfer function always correspond to the whitening on.

The digital one is actually switching. We decided to keep the digital de-whitening on to compensate for the analog one. Otherwise we get a very bad shape of the PDH signal. Sorry Rana...

 I forgot to say that the analog gain of the REFL55 channels has been reduced to 9db

  8376   Fri Mar 29 19:56:02 2013 GabrieleMetaphysicsLSCLock of PRMI on sidebands

I finally managed to get long stretches of PRMI lock, up to many minutes. The lock is not yest very stable, it seems to me that we are limited by some yaw oscillation that I could not trace down. The oscillation is very well visible on POP.

Presently, PRCL is controlled with REFL55_I, while MICH is controlled with AS55_Q. This configuration is maybe not optimal from the point of view of phase noise couplings, but at least it works quite well. I believe that the limit on the length of locks is given by the angular oscillation. I attach to this entry few plots showing some of the lock stretches. The alignment is not optimal, as visible from a quite large TEM01 mode at the dark port.

Here are the parameters I used:

MICH gain -10   PRCL gain -0.1

Normalization of both error signal on POP22_I with factor 0.004

Triggering on POP22: in at 100, out at 20 for both MICH and PRCL.

POP55 demodulation phase -9

MICH and PRCL control signal limits at 2000 counts

 

There is a high frequency (628 Hz) oscillation going on when locked (very annoying on the speakers...), but reducing the gain made the lock less stable. I could go down to MICH=-1.5 and PRCL=-0.02, still being able to acquire the lock. But the oscillation was still there. I suspect that it is not due to the loops, but maybe some resonance of the suspension or payload (violin mode?). There is still some room for fine tuning...

Lock is acquired without problems and maintained for minutes.

Have a nice week-end!

Attachment 1: lock_prmi5.pdf
lock_prmi5.pdf
Attachment 2: lock_prmi6.pdf
lock_prmi6.pdf
Attachment 3: lock_prmi7.pdf
lock_prmi7.pdf
Attachment 4: oscillation.pdf
oscillation.pdf
Attachment 5: lock_prmi8.pdf
lock_prmi8.pdf
  8375   Fri Mar 29 19:23:49 2013 Gabriele, JenneFrogsLSCAnalog whitening filter of REFL55 not switching

We discovered that the analog whitening filter of the REFL55_I board is not switching when we operate the button on the user interface. We checked with the Stanford analyzer that the transfer function always correspond to the whitening on.

The digital one is actually switching. We decided to keep the digital de-whitening on to compensate for the analog one. Otherwise we get a very bad shape of the PDH signal. Sorry Rana...

  8374   Fri Mar 29 17:24:43 2013 JamieUpdateComputersFB RAID power supply replaced

Steve ordered a replacement power supply for the FB JetStor power supply that failed a couple weeks ago.  I just installed it and it looks fine.

  8373   Fri Mar 29 16:35:54 2013 SteveUpdate40m UpgradingEndtable upgrade

 

 My full effort to get the optical table enclosure ready for the lab has failed today.

What I did: cut IR thermashield sheets to size for sides and top and sandwitched them beetween 1" and 1/4" acrylic.

                   The carpenter shop recut the bottom o -ring groove to 0.250" wide and 0.150" deep.

                   O-ring was cut and installed. So this was ready to go lab.

                  NO, I realized that the liner yellow acrylic was not cut correctly. It was larger than 1" wall.

                 The shop is cutting them down to fit and I have to resize IR shields

  8372   Fri Mar 29 16:30:27 2013 SteveUpdate40m Upgrading table leveling mounts go here

Quote:

Quote:

Steve's suggestion for how to level the end table using "swivel leveling mounts":

end-tabel-leveling.png

 

1, level table with 4 swivels, lock nuts  3/16 -vertical alignment    2, lock this position with 4 x 1/4-20 (these are in place already) to hold table in horizontal direction

   Adjustable mount is here to compensate for bad tilting  floor <http://www.mcmaster.com/#6111K52>  

      See grouting plans of the past   http://nodus.ligo.caltech.edu:8080/40m/7248

 

  8371   Fri Mar 29 00:27:06 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : Assembly

I am currently putting together all components so that they are ready to go on the table once leveling and installation of shield box is done. All dirty optics were drag wiped. These are stored in the cupboard along the Y arm.
I could not find the fiber coupling mount on the old endtable. Also the harmonic separator that reflects the trans beam to PDs and camera is labelled Y1-1064 (??) and I don't know what's the deal with this.
I am nearly 70% done with assembling…so the ex-endtable is almost empty.

Yet to do:
1. Mount 2" optics
2. Hunt, gather and mount appropriate lenses

Points I did not notice earlier:
We need some good 2" lens mounts and also order 2" lenses for IPANG and trans beam.

 

IMG_0433.JPGIMG_0434.JPG

  8370   Thu Mar 28 23:06:48 2013 AnnalisaUpdateAuxiliary locking"Alberto" NPRO laser again on PSL table

 "Alberto"NPRO laser has been moved again on PSL table in order to make a measurement of the beat note varying also the PSL temperature.

It is useful because if the PSL temperature would drift  we have to know which is the NPRO temperature that returns the beat.

I'm going to measure it tomorrow.

 

 

  8369   Thu Mar 28 23:00:30 2013 AnnalisaUpdateABSLBeat note of ATF auxiliary laser found

 

The beat note for the ATF lab laser has been found. 

The measurement has been carried out in the same way as described in elog 8368.

The only difference is that in this case I started from a temperature of 35.2 degC, and I reduced it until the minimum which was 30.71 degC. No beat note in this range.

Then I rised on the temperature and I found the first beat note at 41.46 degC. It has been detected at a frequency of about 120 MHz with an RF power of -53 dBm and a frequency fluctuation of about  +/- 5 MHz. 

I tried to improve the alignment to have a stronger beat, but it was the maximum I could reach. Maybe I could increase the power hitting the photodiode, which was 0.453 mW. 

 

 

  8368   Thu Mar 28 19:32:22 2013 AnnalisaSummaryAuxiliary lockingBeat note found!

 

 I plot the variation of the beat note frequency as a function of "Alberto" NPRO laser's temperature.

After some discussion, now I'm going to vary the PSL temperature and find the auxiliary NPRO temperature matching to have the beat note between the two.

Attachment 1: BeatFreq.jpg
BeatFreq.jpg
  8367   Thu Mar 28 12:50:52 2013 JenneUpdateComputersc1lsc is fine

 Manasa told me that she did things in a different order than her old elog. 

She had

(1) ssh'ed to c1lsc and did a remote shutdown / restart,

(2) restarted fb,

(3) restarted the mxstream on c1lsc,

(4) restarted each model individually in some order that I forgot to ask.

However, with the situation as in her "before" screenshot, all that needed to be done was restart the mxstream process on c1lsc. 

Anyhow, when I looked at the OAF model, it was complaining of "no sync", so I restarted the model, and it came back up fine.  All is well again.

  8366   Thu Mar 28 10:44:30 2013 ManasaUpdateComputersc1lsc down

c1lsc was down this morning.

I restarted fb and c1lsc based on elog

Everything but c1oaf came back. I tried to restart c1oaf individually; but it didn't work.

Before:

cds_FE.png

After:

cds_fe1.png

  8365   Thu Mar 28 07:58:15 2013 SteveUpdateGreen LockingNPRO repair

Quote:

 

 JDSU can repair the Lightwave M126-1064-700 NPRO, sn 415  They do not need the Controller sn 516 

 Posted in the 40m Wiki_ PSL_ NPRO  cost repair and/or option to buy Innolight laser as replacement

 NPRO shipped out for evaluation yesterday under RMA 18022707

  8364   Thu Mar 28 04:21:40 2013 ranaUpdateLockingPRMI progress - even better!

 

 A key step was turning off the whitening filters. With the previous setting (G = 15 dB, white on), the error signals (post anti-whitening) had amplitudes of ~500 counts. This means that they can go as high as (150/15)^2 * 500 = 50000 counts on the ADC.

The purpose of the whitening filter is to match the noise / range of the signal to the ADC. What we would like to do is use the minimum gain so as to make the RFPD electronics noise + shot noise be ~equal to the ADC noise. i.e.

sqrt(V_PD^2 + v_shot^2) * G_white = V_ADC

The RFPD noise is ~3 nV before the internal preamp. The MAX4107 has a gain of 10. There is a factor of 1/2 from the voltage division of the RFPD's 50 Ohm series resistor and the input impedance of the mixer. There is also a power splitter between the PD output and the mixer which gives us a 3 dB loss. The mixer has a conversion loss of ~5-6 dB depending upon the LO level.

V_PD = 3e-9 * (10 * 1/2 * 1/sqrt(2) * 1/2) = 5e-9 V/rHz   (this is already bad; the signal coming out of the mixer needs to be amplified by x10 before going out to the whitening board).

In any case, its clear that we need something like 60 dB of gain for the PD noise to match the ADC noise. This is why increasing the whitening gain improves the error signal's SNR, reduces the hash driving the optics, and improves the locking. We should run with 45 dB gain and the switch on whitening after the lock.

 

Even better would be to modify the LT1128 input stage of the card to have the single stage of fixed whitening as we did for iLIGO. Then we can have triple whitening in lock.

  8363   Thu Mar 28 01:59:33 2013 JenneUpdateLockingPRMI progress - even better!

[Gabriele, Rana, Jenne, Jamie, Lisa, Zach]

We tweaked some things after dinner, and our locks got longer (~10sec) and more frequent!

What happened / notes:

* Increased analog gain from 15dB to 27dB for REFL55 I&Q.

* No analog whitening during lock acquisition.  (Need trigger + wait so whitening comes on after ~1sec...but this is not our limitation right now).

* Limit MICH and PRCL control to 5000, so that we don't kick optics too much, which makes them take too long to settle.

* ITMX and ITMY Vio2 filters turned off (PRM still has it on) in the SUS-optic_LSC module.

* MICH and PRCL DoFs triggering on POP22I, with levels 200 & 50.  FMs 4&5 always on.

* MICH and PRCL FM2 triggering on POP22I with levels 400 & 50.

* MICH gain = -0.200

* PRCL gain = +0.150

* MICH and PRCL normalization using POP22I, with matrix values 0.00160 .  This value is ~1/600, where 600 was the peak value of POP22I_ERR.

* REFL 55 phase set back to -15, to minimize PRCL signal in I phase.

* Checked signs for ITMX and ITMY in output matrix for MICH.  Lock MICH using only ITMX or ITMY, find sign to hold on the dark fringe for each.  +1 for ITMY, -1 for ITMX was correct.

* Tweaked up the oplev servos.  See separate elog 8362.  May need more tweaking, such as increasing the UGF, engaging 1Hz resonant gain.

* May need better coil actuator balancing on suspensions at 1Hz.

* Found a weird thing in DTT, which went away after closing and reopening, when looking at time series.  Sometimes we would see a square wave-like jump in the signals, all signals at the same time, with a frequency of 16.6Hz.  This was not present in other data retreival programs, like Jamie's getdata python script.

* We are not sure right now why we are falling out of lock.  We need to investigate more signals, to try to figure out what our current problem is.

* Reduced the amount of misalignment with the "misalign" script, to reduce hysteresis.

To Do / ideas:

* Calibrate oplev signals - see if one optic is moving more than others.

* Calibrate ERR and CTRL  - look at CTRL in meters, see if cavities are moving around like crazy.

* Calibrate POP22 using something like an AM laser modulation trick into units of PRCL SB gain. Compare with expectation - are we locked optimally, or do we have more power that we can be getting out?

* Try feeding back PRCL CTRL to MC2, to make the laser to follow the power recycling cavity, in hopes of reducing angular motion.  Rana tried this quickly with a 1 in the output matrix, but this kicked the MC out of lock - need to try smaller values.

 

  8362   Thu Mar 28 00:31:11 2013 JenneUpdateSUSPRMI optics' oplev servos tuned

[Rana, Gabriele, Jenne, Jamie, Lisa, Rana]

We have tuned the oplev servos for PRM, BS, ITMX, ITMY.  For each, we measured the servo transfer function.  Most had a UGF ~ 3Hz.  For those, we increased the gain by a factor of 2, and engaged the 3.3Hz resonant gains. The other case, such as PRM yaw, the gain was already okay, we just needed to engage the resonant gain.  We also checked the new phase margin, and for some of them switched the elliptic lowpass to 50Hz rather than 30 or 35.

Before and afters:

PRM_OLPIT_tuning_27Mar2013_RedIsAfter.pdf

PRM_OLYAW_tuning_27Mar2013_RedIsAfter.pdf

bs_pit.pdf

bs_yaw.pdf

ITMY_OLPIT_tuning_27Mar2013_RedIsAfter.pdf

ITMY_OLYAW_tuning_27Mar2013_RedIsAfter.pdf

itmx_pit1.pdf

itmx_yaw.pdf

We need to, as a last check, look at the spectra before and after to ensure that no modes (like bounce or roll) are newly excited.

  8361   Wed Mar 27 21:53:21 2013 AnnalisaUpdateABSLBeat note of ATF auxiliary laser

After measuring the beat note, the "Alberto" NPRO auxiliary laser has been moved from the PSL table to the POY table. Its beam profile is going to be measured. It's going to be used as green laser on the END table, in place of the broken one.

The auxiliary laser borrowed form ATF lab (which will be used for the ABSL measurement) has been set on the PSL table to make a measurement of the beat note between it and the main laser.

The setup is mostly the same of the previous beat note measurement . In this case, laser input power is 326 mW, so I needed to replace one of the mirrors of the steering optics with a BS 50% reflecting in order to have less than 1 mW on the PD.

Now, the total power on the PD is less than 0.5 mW.

I didn't measure the beat note yet to leave the PSL table as quite as possible for the locking procedures.

To do:

Measure the beat note, fiber coupling the NPRO laser to bring it to the POY table.

 

  8360   Wed Mar 27 15:46:13 2013 JenneUpdateLockingPRMI progress

[Gabriele, Jenne]

We have achieved PRMI locks of the order ~5 seconds! Here is an example lock:

PRMI_sb_27Mar2013.png

This was actuating MICH on the ITMs (+1 for ITMY, -1 for ITMX in the output matrix), and PRCL on PRM (+1). 

PRCL gain was +1, MICH gain was -10.

PRCL signal was normalized with POP22I with a matrix value of 0.003 . (No normalization of MICH).

Both PRCL and MICH were triggered on POP22I with high thresh of 200, low of 50.  MICH and PRCL FM2 (integrators) were triggered on POP22I with thresh of 400 and low thresh of 50.  FMs 4 and 5 were on for both MICH and PRCL always.

We zoomed in on the MICH_OUT signal, and the instability looks like it is around 300 Hz.  We aren't sure what this is.  I think this is a similar frequency to an oscillation that Yuta saw, but I'll have to check the old elogs.

PRM and BS SUS_LSC_POS filter banks both have notches between 1280-1290Hz.  The ITMs do not have this "Vio2" filter.

This is one of the first locks with the new triggering of both the INPUT and OUTPUT of the control filter banks.  I modified the lsc model before lunch. 

To do:  Where is this 300Hz coming from, and what can we do about it?  Why are we losing lock?  It's not due to the oscillation - maybe too much afternoon seismic?  Steve says he went next door and the rock monster / river is on medium/high.

 

  8359   Tue Mar 26 20:20:10 2013 CharlesUpdateISSISS Design Plans - Servo Noise Analysis

In order to allow other individuals besides myself to consider the proposed design of the ISS, I have created a publicly available CircuitLab drawing, which can be found here: CircuitLab Drawing. For simplicity, I have used ideal op-amps without voltage rails or their associated power supplies. In the actual implementation of the ISS, we will most likely also have trim resistors to ensure a zero offset for each op-amp. We interpret the PD as a voltage source for simplicity and I will use an actual summing amplifier in place of the summing junction used in the diagram.

The diagram linked above is simply a naive copy of a design by Rich Abbott so there are most likely mistakes and/or unnecessary elements, but it is a work in progress. I began discussing, with Jamie, the relative use of the first few filter stages in the servo. As far as my understanding goes, the first 'stage' was part of cascade of op-amps that served to convert a differential input from the PD into a single DC signal referenced to ground. Indeed, the first stage of my diagram (U1) is simply a unity-gain low-pass filter with f~5 MHz. Additionally, the second filter 'stage', U2, is also a unity-gain low-pass filter although it introduces a phase shift of 180 deg as the input to the second stage is on the inverting input of the op-amp. These characteristics were determined using LISO and examining the transfer function.

Noise analysis was also performed for the above circuit. The noise from various elements is examined at the output of the servo (labeled as 'outU6' in my LISO file). In the attached diagram, we see the voltage noise at the output from each op-amp as well as the sum of all the various noises, which includes resistor noise and current noise from the inputs of each op-amp. These are LISO's standard considerations and it is also worthwhile to note that the result is not referred to the circuit input, but as we have the transfer function of the whole servo, referring the noise to the input is trivial.

I have also included the following output for the sake of completeness.

from 1 Hz onwards noise by OP:I+ (U3) dominates.

from 38.6812 Hz onwards noise by R(R24) dominates.

from 115.478 Hz onwards noise by R(R11) dominates.

 

 

Attachment 1: ISS.pdf
ISS.pdf
  8358   Tue Mar 26 17:32:30 2013 ChloeUpdate QPD/ECDL Progress

I built the summing/subtracting circuit on the breadboard, and hooked this up with one of the other QPDs I found (image of setup attached). I wasn't able to get this to read the correct signals when testing with a laser pointer after a couple of hours of troubleshooting... I will hopefully get this working in the next day or 2...

I'm going to read up on ECDL stuff for Tara tonight and hopefully figure out what sort of laser diode we should purchase, since I'm meeting with Tara tomorrow. experimenting

Attachment 1: IMG-20130326-00245.jpg
IMG-20130326-00245.jpg
  8357   Tue Mar 26 17:27:17 2013 KojiUpdateIOOTuning MC length/FSR

Please add the discussions on the cavity absolute length (and its change, adjustment precision),
identification of the peaks, before/after comparison of the plot, the effect of the MC REFL PD response,
and comparison of the cavity linewidth vs deviation of the 55MHz SBs from the resonance.

  8356   Tue Mar 26 16:36:24 2013 SteveUpdate40m Upgrading table leveling mounts go here

Quote:

Steve's suggestion for how to level the end table using "swivel leveling mounts":

end-tabel-leveling.png

 

1, level table with 4 swivels, lock nuts  3/16 -vertical alignment    2, lock this position with 4 x 1/4-20 (these are in place already) to hold table in horizontal direction

   Adjustable mount is here to compensate for bad tilting  floor <http://www.mcmaster.com/#6111K52>  

      See grouting plans of the past   http://nodus.ligo.caltech.edu:8080/40m/7248

Attachment 1: optabSTAND.PDF
optabSTAND.PDF
Attachment 2: swivvelMountEE.jpg
swivvelMountEE.jpg
  8355   Tue Mar 26 16:10:31 2013 JamieUpdate40m UpgradingETMY table leveling

Steve's suggestion for how to level the end table using "swivel leveling mounts":

end-tabel-leveling.png

 

  8354   Tue Mar 26 15:43:45 2013 ranaUpdateIOOTuning MC length/FSR

  Please attach the data file of the measurement - its hard to get the real information out of picture. In general, WE should always include the code / explanation of how to reconstruct the plot and get the scientific information out.

  8353   Tue Mar 26 15:00:55 2013 ManasaUpdateIOOTuning MC length/FSR

Quote:

[Koji, Manasa]

We looked at the MC modulation frequency on the spectrum analyzer and observed beat notes between MC modulation freq (29.5MHz) and modulation frequencies (11.06 MHz and 55.3MHz).

Beat notes were suppressed by changing the carrier frequency from 11.065910 MHz to 11.066147MHz. 

Detailed discussion and data will be posted in the next elog.

The goal was to tune the carrier modulation frequency, f1 ~ 11.06 MHz to match the FSR  (c/2L) of the MC. (Reference to the technique: R.G.DeVoe et al., PRA 30, 5, Nov 1984)

We looked at the MC_REFL output on the spectrum analyzer. Since the MC FSR was not well matched with the carrier modulation frequency, we observed significant beat notes at the following frequencies (fMC-f1),  (fMC+f1), (fMC-f2) and (fMC+f2); where fMC (the MC modulation frequency) = 29.5MHz, f1(carrier modulation frequency) ~ 11.06MHz and f2 ~ 55.3MHz. The carrier modulation frequency was changed at the frequency generator until these beat notes were suppressed i.e. until the cavity FSR matched the carrier modulation frequency.

The plot below shows the MC spectrum after the beat notes were suppressed.

MC_spectrum.png

Attachment 2: MC_spectrum.zip
  8352   Tue Mar 26 11:33:55 2013 JamieUpdateOpticsHOWTO calculate effective RoC of flipped TT

In case anyone is curious how I got the numbers for the effective radius of curvature of the flipped TT mirrors, I include the code below.  Now you can calculate at home!

Here's the calculation for the effective RoC of a flipped SR2 with nominal un-flipped HR RoC of -600:

>> [Mt, Ms] = TTflipped(600, 5);
>> M2Reff('t', Mt, 5)

ans =

  412.9652

>>

Attachment 1: mirror.m
function [Mt, Ms] = mirror(R, a1, n)
% [Mt, Ms] = mirror(R, a1, n)

if length(R) > 1
    Rt = R(1);
    Rs = R(2);
else
    Rt = R;
    Rs = R;
end
... 9 more lines ...
Attachment 2: transmission.m
function [Mt, Ms] = transmission(R, a1, n1, n2)
% [Mt, Ms] = transmission(R, a1, n1, n2)

if length(R) > 1
    Rt = R(1);
    Rs = R(2);
else
    Rt = R;
    Rs = R;
end
... 19 more lines ...
Attachment 3: TTflipped.m
function [Mt, Ms] = TTflipped(R, a1)
% [Mt, Ms] = TTflipped(R, a1)

if length(R) > 1
    Rt = R(1);
    Rs = R(2);
else
    Rt = R;
    Rs = R;
end
... 32 more lines ...
Attachment 4: M2Reff.m
function Reff = M2Reff(type, M, a)
% Reff = M2Reff('type', M, a)

n = 1;

R = -2*n/M(2,1);

ca = cos(a*pi/180);

switch lower(type)
... 8 more lines ...
  8351   Tue Mar 26 11:08:03 2013 JenneUpdateLockingPRMI locking should be possible - let's get it done!

[Jenne, Gabriele, Jamie]

We have looked at Koji's old Finesse code, and determined that the PRMI sensing matrix that he calculated was for the sideband-resonant case.  Thus, this is the sensing matrix we are interested in for locking. Gabriele has confirmed this using independently written code with his software, Mist.

Pics or it didn't happen:

 PRMI_SBres_plot_allTheFs_LowRes.png

The sensing matrix is:

LSC_sigmatrix_PRMI.pdf

Numerical values are on the wiki:  https://wiki-40m.ligo.caltech.edu/IFO_Modeling/SensingMatrix

For any of the REFL PDs (1*f1, 1*f2, 3*f1, 3*f2), the PRCL signal is a factor of ~100 larger than the MICH signal.  Rana assures us that, with some clever triggering, we should be able to lock the PRMI. 

This means that we will not be venting in the next few days to flip the SRC folding mirror.  We will work on PRMI lock (probably first with 1*f, but then quickly moving on to 3*f), and as soon as Annalisa and Manasa have the new ETMY table ready for us, we will then do PRFPMI.  (We can also play with the Xarm green until Yarm green is back).

  8350   Tue Mar 26 08:50:02 2013 SteveUpdateGeneraljanitor day

Keven, our janitor accidentally pushed the main entry door laser emergency stop switch.

The laser was turned back on. The MC and the arms were  started  flashing happily as they were before.

  8349   Tue Mar 26 01:40:49 2013 KojiUpdateIOOTuning MC length/FSR

I'm still waiting for the follow-up analysis of the modulation freq tuning.

  8348   Tue Mar 26 00:17:47 2013 JenneUpdateLockingREFL pickoff fraction

To see how much of the light that comes out of the REFL port actually goes to the PDs, I measured the power immediately after leaving the vacuum (~575mW) and in front of REFL11 (~5mW) and REFL55 (~6mW).

So, 0.01 of the power leaving the vacuum actually goes to the REFL PDs. This number will be useful when calculating the actual signals (in volts) that we expect to see.

  8347   Tue Mar 26 00:06:37 2013 JenneUpdateLockingAS beam put back on PD

[Jenne, Gabriele]

We aligned MICH (first locked Yarm, but didn't optimize since we don't have TRY, then locked Xarm, then aligned MICH), but there was no beam on AS55.  We went out to check, and the beam was almost not hitting the small steering mirror between AS55.  We adjusted the BS splitting the beam between camera and PD, and got the beam back on AS55. We could then lock MICH.

We also futzed with the REFL55 phase to get PRCL stuff in I, and MICH stuff in Q.  The procedure was to align PRMI, then kick PRM in pos, and adjust the phase so we got signal mostly in I after the kick.  We started at the original value of 60deg, but are leaving it at -20deg.

  8346   Mon Mar 25 23:27:26 2013 ChloeUpdate QPD Circuit

In order to test the mount vibrations, I will likely try and make a different circuit work (with the summing/subtracting on an external breadboard) and designing an optimal circuit will be a side project. This is the circuit with the power supply Rana came up with, and the design I had in mind for the rest of the circuit. In my free time, I will try to figure out what parts to get that reduce noise and slowly work on building this, since it would be useful to have in the lab. 

 https://www.circuitlab.com/circuit/7sx995/qpd-circuit/#postsave_access_control_settings

  8345   Mon Mar 25 23:20:57 2013 AnnalisaSummaryAuxiliary lockingBeat note found!

[Annalisa, Manasa]

The beat note between the main PSL and the auxiliarly NPRO has been found!

The setup didn't change with respect to the one described on the previous note on the elog. A multimeter has been connected to the laser controller diagnostic pin to read out the voltage that indicated the laser crystal temperature.

The connector has been taken from the Yend table laser controller.

The voltage on the multimeter gave the same temperature shown by "Laser temperature" on the display of the controller, while "set temperature" was wrong.

The temperature has been varied using the laser controller with reference to the voltage read on the multimeter display.

Starting from 35.2 °C, the temperature has been first lowered until 20 °C and no beat note has been found, then temperature has been increased up to 35.2 °C and the first beat note has been found at 38.0 °C.

It has been detected at a frequency of about 80 MHz with an RF power of -27 dBm and a frequency fluctuation of about  +/- 4 MHz.

To do:

I made more measurements slowly varying the laser temperature, to see how the beat note frequency changes with it. I'll make the plot and post it as soon.

  8343   Mon Mar 25 23:17:11 2013 ranaSummaryIOOMC

I measured the dark noise and regular intensity noise in MC trans tonight to get some estimate for the free running spectrum that the Chas ISS must overcome. PDF is attached.

XML file is /users/rana/dtt/MC-trans_130325.xml

The RIN normalization was done by using the mean of the SUM time series. The dark noise was measured by misaligning MC2 and making the same measurement with the bright normalization.

Attachment 1: MC-trans.pdf
MC-trans.pdf
  8342   Mon Mar 25 18:58:09 2013 Albert YangUpdateoptical tablesOptical Table Toolboxes

For those of you who spend annoying amounts of time looking for tools, fear no more. Toolboxes for each optical table are coming!

They will probably have:

IR Viewer (a few optical tables will have IR viewers, these specific tables will be labeled in the diagram coming out later)

IR Card

Ball screw drivers (3/16 in.) 6-8 in. handle

SMA Wrench

Allen Keys

Flashlight

Various Connectors (I'll find out what's needed at some point)

Zip Ties

Small flat screwdrivers (for adjusting camera gains)

 

Please suggest what else may be needed in these boxes.

The boxes will be held to the side of the tables, either by magnets or screws. A diagram of where they will be placed on each optical table in order to minimize obstruction of walkways will be distributed soon. Any objections can then be noted.

  8341   Mon Mar 25 18:42:38 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : Timeline

 Endtable upgrade - timeline and progress chart. End table upgrade wiki page updated.

 

Attachment 1: Endtable_progress.pdf
Endtable_progress.pdf
  8340   Mon Mar 25 18:07:31 2013 SteveUpdate40m UpgradingETMY table is out

Quote:

Quote:

Yes! We are swapping.

I'll be there very soon!

 The ETMY optical table 4' x 2' with all optical components was placed on two carts and rolled out of the end.  The 4' x 3' x 4" TMC 78-235-02R breadboard was placed into position and marked for anchoring bolt locations.

It will be drilled, tapped, shimmed, leveled and bolted it into final position tomorrow morning.  I'm planning to bring  the acrylic enclosure to the east end tomorrow afternoon.

 The new table is tapped and leveled, but not ready for final anchoring. The existing 8 mm shims on the top of the support legs will be moved to the bottom of the legs, where more torque is available on the 1/2" bolts

The coated tin film sheets are being installed at the shop. The table and the enclosure will be ready on Friday.

Attachment 1: shimmingEE.jpg
shimmingEE.jpg
  8339   Mon Mar 25 16:51:59 2013 JenneUpdateOpticsupdated calculations of PRC/SRC g-factors and ARM mode matching

I think we like the idea of flipping SR2 better than SR3 for the same ghost beam reasons as PR2 is better than PR3.  There isn't very much free space in the BS chamber, and if we flip PR/SR3, we have to deal with green ghosts as well as IR.

The flipping SR2 case seems okay to me - flipping either one of the SR folding mirrors gives us a slightly better g-factor than the design with infinite curvatures, and flipping SR2 gives us slightly better mode matching to the arm than the flipped SR3 case, but more importantly, there are fewer ghosts to deal with.  I vote we flip SR2, not SR3.

  8338   Mon Mar 25 16:51:43 2013 ChloeUpdate Final QPD Circuit Design

This is the final version of the QPD circuit I'm going to build. After playing around with the spatial arrangement, this should fit into the box that I was planning to use, although it will be a rather tight fit. The pitch, yaw, and summing circuit will be handled with a quad op amp. Planning to meet with Eric tomorrow to figure out the logistics of building things.

In the meantime, I'm reading about designing the ECDL for my summer project with Tara. He sent me several papers to read so we can talk on Wednesday.

Attachment 1: IMG-20130325-00244.jpg
IMG-20130325-00244.jpg
  8337   Mon Mar 25 15:24:41 2013 ranaUpdateLockingTT1 drift

 

 I doubt that it has truly disappeared. Can you please make a measurement of it and quantify the hysteresis using some angle sensor?

  8336   Mon Mar 25 15:01:33 2013 ManasaUpdateLockingTT1 drift

Quote:

2. TT1 drifting in pitch (Bistable)
During the arm alignment routine for spot centering, we observed that TRY dropped (from TRY = 0.9 until the arm lost lock) every 40minutes or so. The arm was relocked by moving TT1 in pitch. The (locking - unlocking due to drift - relocking) cycle was monitored and we observed that it was bistable i.e. if TT1 was moved up in pitch (0.2 on the slider) to relock for the first time ; the next time it lost lock, TT1 had to be moved down by nearly the same distance to relock the arm.
Moving TT2 or the testmasses did not help with relocking the arms; so TT1 seems to be the one causing all th trouble atleast for today.

Update

TT1 was moved in pitch to restore flashing in the arms on Wednesday (Mar 20) so that it doesn't drift too far off making it difficult to lock again. Since then, the arms have been flashing without any drifting and TTs have remained untouched. The hysteresis has disappeared.

  8335   Mon Mar 25 11:42:45 2013 JamieUpdateComputersc1lsc mx_stream ok

I'm not exactly sure what the problem was here, but I think it had to do with a stuck mx_stream process that wasn't being killed properly.  I manually killed the process and it seemed to come up fine after that.  The regular restart mechanisms should work now.

No idea what caused the process to hang in the first place, although I know the newer RCG (2.6) is supposed to address some of these mx_stream issues.

  8334   Mon Mar 25 09:52:22 2013 JenneUpdateComputersc1lsc mxstream won't restart

Most of the front ends' mx streams weren't running, so I did the old mxstreamrestart on all machines (see elog 6574....the dmesg on c1lsc right now, at the top, has similar messages).  Usually this mxstream restart works flawlessly, but today c1lsc isn't working.  Usually to the right side of the terminal window I get an [ok] when things work.  For the lsc machine today, I get [!!] instead. 

After having learned from recent lessons, I'm waiting to hear from Jamie.

  8333   Fri Mar 22 23:23:38 2013 AnnalisaUpdateLockingBeat note still missing

[Annalisa, Manasa, Koji]

I updated the setup for the beat note. The main reason is that I needed to keep the ADJ to 0 in way to operate at the nominal laser power.

Now the input power of the laser is increased (about 315 mW) and needs to be dumped so as not to exceed the PD threshold of 1mW. 

Moreover, a lens has been added to match the two beams size.

A BS has been removed from the PSL pick-off beam path, so the PSL power hitting the BS is now about 100 uW, and the total power on the PD is 0.7mW.

I also verified that both the beams are S polarized.

To find the beat note, the laser temperature has been varied  through the laser controller and not adding a Voltage with the power supply.

A range of temperature of 30 degC has been spanned, but we suppose there should be some calibration problem with the controller, since set temperature is not the same as Laser temperature on the display. 

Anyway, no beat note has been found up to now.

 

An external monitor has to be added to check the real temperature of the crystal.

The next possible plan is to vary the PSL temperature and try to find the beat note. 

 

P.S.: The HEWELETT PACKARD 8591E spectrum analyzer works! The monitor only took some time to turn on!

 

Attachment 1: Beat_note_-_new_setup.jpg
Beat_note_-_new_setup.jpg
  8332   Fri Mar 22 19:46:29 2013 KojiSummaryLSCDiode impedance test result

I've tested Perkin-Elmer InGaAs PDs at OMC Lab.

- The diode impedances were measured with the impedance measurement kit. Reverse bias of 5V was used.

- Diode characteristics were measured between 10MHz and 100MHz.

- 4-digit numbers are SN marked on the can

- Ls and Rs are the series inductance and resistance

- Cd is the junction capacitance.

- i.e. Series LCR circuit o--[Cd]--[Ls]--[Rs]--o

C30665GH, Ls ~ 1nH

0782 Perkin-Elmer, Rs=8.3Ohm, Cd=219.9pF
1139 Perkin-Elmer, Rs=9.9Ohm, Cd=214.3pF
0793 Perkin-Elmer, Rs=8.5Ohm, Cd=212.8pF

C30642G, Ls ~ 12nH

2484 EG&G, Rs=12.0Ohm, Cd=99.1pF
2487 EG&G, Rs=14.2Ohm, Cd=109.1pF
2475 EG&G glass crack, Rs=13.5Ohm, Cd=91.6pF
6367 ?, Rs=9.99Ohm, Cd=134.7pF
1559 Perkin-Elmer, Rs=8.37Ohm, Cd=94.5pF
1564 Perkin-Elmer, Rs=7.73Ohm, Cd=94.5pF
1565 Perkin-Elmer, Rs=8.22Ohm, Cd=95.6pF
1566 Perkin-Elmer, Rs=8.25Ohm, Cd=94.9pF
1568 Perkin-Elmer, Rs=7.83Ohm, Cd=94.9pF
1575 Perkin-Elmer, Rs=8.32Ohm, Cd=100.5pF

C30641GH, Perkin Elmer, Ls ~ 12nH

8983 Perkin-Elmer, Rs=8.19Ohm, Cd=25.8pF
8984 Perkin-Elmer, Rs=8.39Ohm, Cd=25.7pF
8985 Perkin-Elmer, Rs=8.60Ohm, Cd=25.2pF
8996 Perkin-Elmer, Rs=8.02Ohm, Cd=25.7pF
8997 Perkin-Elmer, Rs=8.35Ohm, Cd=25.8pF
8998 Perkin-Elmer, Rs=7.89Ohm, Cd=25.5pF
9000 Perkin-Elmer, Rs=8.17Ohm, Cd=25.7pF

 

Note:  Calculated Ls&Rs of straight wires
  1mm Au wire with dia. 10um -> 1nH, 0.3 Ohm
20mm BeCu wire with dia. 460um -> 18nH, 0.01 Ohm

  8331   Fri Mar 22 01:28:56 2013 ManasaUpdateLasersBeam profile of NPRO from ATF

The NPRO from ATF has been installed on the POY table.

I have been making measurements to characterize the beam profile of this laser. I am using an AR coated laser window as a beam sampler at 45deg and the razor blade technique to measure the beam size along z. Details of the procedure along with analysis and results from this will follow.

  8330   Thu Mar 21 17:06:34 2013 SteveUpdate40m UpgradingETMY table is out

Quote:

Yes! We are swapping.

I'll be there very soon!

 The ETMY optical table 4' x 2' with all optical components was placed on two carts and rolled out of the end.  The 4' x 3' x 4" TMC 78-235-02R breadboard was placed into position and marked for anchoring bolt locations.

It will be drilled, tapped, shimmed, leveled and bolted it into final position tomorrow morning.  I'm planning to bring  the acrylic enclosure to the east end tomorrow afternoon.

Attachment 1: ETMYtableSupport.jpg
ETMYtableSupport.jpg
  8329   Thu Mar 21 14:02:45 2013 AnnalisaUpdateLockingincident angles

Quote:

Is there a reason to use non-45 degree incident angles on the steering mirrors between the laser and the PD?  I would always use 45 degree incident angles unless there is a really good reason not to.

 Actually not, it is a mistake! It is one of the things I'm going to modify, in addition to add more BS to reduce the power. 

  8328   Thu Mar 21 13:50:08 2013 JamieUpdateLockingincident angles

Is there a reason to use non-45 degree incident angles on the steering mirrors between the laser and the PD?  I would always use 45 degree incident angles unless there is a really good reason not to.

  8327   Thu Mar 21 13:11:42 2013 AnnalisaUpdateLockingFinding the beat note

Quote:

Give us more info on the elog:
What PD are you using? How much power the beams on the recombining BS are? What kind of BS is it?
How are you looking for the beat note? (on the scope? or spectrum analyzer?)
What was the scanned temp range?

Three points to be checked:

- Polarization

- Alignment

- Temperature

Experimental Setup

I'm using a 1611 New Focus PD (1 GHZ, with maximum input power 1mW), and the total power hitting on the PD is of about 0.650 mW.

The current of the NPRO laser is set to 1.38 A, so that the input power is 19 mW. The beam is initially damped by a 10% reflection BS and then it hits a 33% reflection BS (where it recombines with the PSL pick-off beam) with 2 mW power.

After this second BS the power is reduced to 0.592 mW.

The PSL pick-off hits on the 33% reflection BS with 65.5 uW power, and it exit with a 47 uW power.

I connected a power supply to apply a Voltage to the slow frequency BNC, in way to tune the laser frequency.

I'm using the AGILENT 4395A Spectrum analyzer to make the measurement. I tried to use the HEWELETT PACKARD 8591E spectrum analyzer, but the monitor didn't turn on.

The temperature spanned until now in only of about 10 deg C, because I realized that I needed a better alignment, so I added a lens in front of the PD and I did a better alignment. 

Moreover, the current of the laser is too low, so I need to increase it and add more beam splitters in the beam path to dump the beam, in way to don't reach the PD threshold.

I knew that both the beams are s-polarized, but maybe I can check it again.

Attachment 1: Beat_note_setup.jpg
Beat_note_setup.jpg
ELOG V3.1.3-